OpenStudy (anonymous):

CAN SOMEONE CHECK MY ANSWERS PLEASE???? 1.Which is the converse of the statement? If x is odd, then 2x + 1 is odd. If 2x + 1 is odd, then x is odd. If 2x + 1 is not odd, then x is not odd.<<<< If x is odd, then 2x + 1 is odd. If x is not odd, then 2x + 1 is not odd. 2.Simplify. -27a^4b^3/3ab –9a2b –9a3b2 –3ab -a^2/9b<<<<<<<<<<<

OpenStudy (anonymous):

@UsukiDoll

OpenStudy (usukidoll):

ok so the negation of the first statement is If x is NOT odd, then 2x + 1 is NOT odd. converse is flipping the statement If 2x+1 is odd, then x is odd contrapositive is the negative version of that converse statement if 2x+1 is NOT odd, then x is NOT odd

OpenStudy (usukidoll):

\[\frac{-27a^4b^3}{3ab} \rightarrow = -9a^3b^2\]

OpenStudy (usukidoll):

exponent laws \[\frac{a^4}{a} \rightarrow a^{4-1} \rightarrow a^3 \] same applies to b

OpenStudy (anonymous):

Thats for the second question right?

OpenStudy (usukidoll):

yeah... imo both of your answers are way off

OpenStudy (anonymous):

Oh..

OpenStudy (usukidoll):

Original statement If A then B Negation If NOT A then NOT B conversely if B then A contrapositive if NOT B then NOT A

OpenStudy (usukidoll):

Original statement A = x is odd B = 2x+1 is odd If A then B if x is odd then 2x+1 is odd Negation If NOT A then NOT B if x is NOT odd then 2x+1 is NOT odd conversely if B then A If 2x+1 is odd then x is odd contrapositive if NOT B then NOT A if 2x+1 is NOT odd then x is NOT odd

OpenStudy (usukidoll):

these guys fail to mention that you can fine tune the negative versions of the statement If NOT A then NOT B if x is NOT odd then 2x+1 is NOT odd if x is even then 2x+1 is even

OpenStudy (anonymous):

isnt it odd?

OpenStudy (usukidoll):

I said you can refine the negative versions of the statement Negation If NOT A then NOT B if x is NOT odd then 2x+1 is NOT odd if it's not odd, then it's even.. but that can be easily disproved because 2k is even and 2k+1 is odd

OpenStudy (anonymous):

Alright..I understood what you told me and when I used it in the equation I get a different answer than they give in the options

OpenStudy (usukidoll):

a different answer?

OpenStudy (anonymous):

Yes I get an even but the choices are all odd

OpenStudy (usukidoll):

it's probably because they haven't realized that they can fine tune it .

OpenStudy (usukidoll):

but converse statements are usually positive ... it's the original if A then B statement written in reverse.

OpenStudy (anonymous):

So C should be the reverse from the original A then B

OpenStudy (anonymous):

Am I correct? or still way off?

OpenStudy (usukidoll):

worse.......and on top of that I gave out the definitions earlier

OpenStudy (usukidoll):

look at what I wrote earlier.

OpenStudy (anonymous):

So if 2x+1 is odd, then x is odd right?

OpenStudy (usukidoll):

yes

OpenStudy (anonymous):

So if I got that then the answer is A

OpenStudy (usukidoll):

yes

OpenStudy (anonymous):

YES!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

Can you help with the first question?

OpenStudy (usukidoll):

of?

OpenStudy (anonymous):

Sorry I meant the second question

OpenStudy (usukidoll):

it's straight forward..... you need to understand the exponent laws for a and b we know that \[\frac{-27}{3} = -9\]

OpenStudy (usukidoll):

if a and b are just single then it's technically \[a^1\] and \[b^1\] but no one writes it like that anymore

OpenStudy (usukidoll):

the same goes for -1c... professors will mark points off not because it's wrong.. because the notation is childish... so the correct way to express your answer is -c

OpenStudy (anonymous):

Oh really? how come?

OpenStudy (usukidoll):

they claim that you grew up from high school so now it's time to write it in college style notation

OpenStudy (usukidoll):

the same goes for simple stuff like this 7+(-7) you can't have a double sign it's 7-7 =0 again for 7-(-7). distribute that negative it's now 7+7 = 14

OpenStudy (usukidoll):

when I reached Algebra 2, I had to knock it off with writing -1c and just write -c for the simple problems, it's Algebra 1 that you need to stop writing double signs

OpenStudy (usukidoll):

and stop writing a^1

OpenStudy (usukidoll):

that's how the professors work now.

OpenStudy (anonymous):

You dont know how many questions I got wrong b/c this -3-3 is actually -3+3 or 5-5-4 is actually 5+ -5 + -4

OpenStudy (usukidoll):

you can't have double signs... can't have +-, -- , -+,++ . just one is allowed either - or +

OpenStudy (usukidoll):

I don't know why teachers write stuff like that...it's irritating

OpenStudy (anonymous):

That's how I've had it for a few tests

OpenStudy (anonymous):

It is reasons why I come to open study xD

OpenStudy (usukidoll):

what grade are you in?

OpenStudy (anonymous):

I'm a freshman

OpenStudy (usukidoll):

oh 9th grade?

OpenStudy (anonymous):

Yeah

OpenStudy (usukidoll):

where did you get that first question? It belongs in logic and discrete math. something in the 300 level at my university

OpenStudy (usukidoll):

unless you're at a private school?

OpenStudy (anonymous):

No I'm home schooled

OpenStudy (usukidoll):

oh O_O... in a way that's great because public schools went nuts... you don't want to learn common core..trust me. I'm a Math Major and I hate everyday common core did to Math .

OpenStudy (usukidoll):

*everything

OpenStudy (anonymous):

Lmao if you dont like it then why major in it?

OpenStudy (usukidoll):

If I didn't go further in my Math Studies, I wouldn't have been able to help you.

OpenStudy (usukidoll):

xD. I said Common Core Math was awful....the stuff that makes simple subtraction crazy. my cousins, second grade and kindergarten, are stuck with it. Public and private schools have it. x(